### v1

# Relativistic Axion with Nonrelativistic Momenta: A Robust Bound on Minimal ALP Dark Matter

Abstract
The axion‐like particle (ALP), a pseudo Nambu-Goldstone boson that couples to two photons, has been studied extensively in recent years as a dark matter candidate. For initial field configurations in a minimal ALP model explaining the observed dark matter abundance, we need the potential height to exceed the ALP energy density at redshift $z\approx 5.5\times 10^{6}$ leading to:
$f_{\phi}\gtrsim 4\times 10^{13}\,\mathrm{GeV}\,\biggl{(}\frac{10^{-18}\,\mathrm{eV}}{m_{\phi}}\biggr{)},$
where $m_{\phi}$ and $f_{\phi}$ denote the ALP mass and decay constant, respectively. This bound is known for the ALP dark matter dominated by the homogeneous zero‐momentum mode, under the requirement that coherent oscillations begin early enough to satisfy the late‐forming dark matter constraint. One loop hole to evade this limit may be to introduce a large amount of the non-relativistic modes of the ALP with non-vanishing momenta. Here we show that the same limit remains valid even if nonzero‐momentum modes dominate. Interstingly, when nonrelativistic gradient modes prevail, the ALP behaves relativistic radiation rather than matter, if it violates the limit. Moreover, if the typical momentum is sufficiently small, domain walls form.
I Introduction
The axion or axion-like particle (ALP) is a pseudo Nambu-Goldstone boson whose mass arises from the explicit breaking of its shift symmetry (see Refs. Jaeckel and Ringwald (2010); Ringwald (2012); Arias et al. (2012); Graham et al. (2015); Marsh (2016); Irastorza and Redondo (2018); Di Luzio et al. (2020) for reviews). A well-known example is the QCD axion, which provides a solution to the strong CP problem via the Peccei-Quinn mechanism Peccei and Quinn (1977a, b); Weinberg (1978); Wilczek (1978).
Dark matter provides direct evidence for physics beyond the Standard Model, yet its nature remains unknown. Among the proposed candidates, the axion is particularly well motivated. In the misalignment mechanism Preskill et al. (1983); Abbott and Sikivie (1983); Dine and Fischler (1983), the axion begins coherent oscillations when its mass becomes comparable to the Hubble parameter, and the resulting energy density can account for dark matter. It has also been shown that if inflation lasts sufficiently long and the Hubble parameter during inflation is lower than the QCD scale, the predicted axion abundance can be significantly altered Graham and Scherlis (2018); Takahashi et al. (2018); Ho et al. (2019); Alonso-Álvarez et al. (2020). Additionally, the axion/ALP can be initially placed near the hilltop of its potential through interactions with other fields, thereby delaying the onset of oscillations and enhancing its abundance Daido et al. (2017); Co et al. (2019); Takahashi and Yin (2019); Nakagawa et al. (2020); Narita et al. (2023). Light ALPs can also be produced via stimulated emission, such as from inflaton, Higgs decay or even thermal scattering Moroi and Yin (2021a, b); Nakayama and Yin (2021); Yin (2023); Sakurai and Yin (2025), or through wide resonance effects Co et al. (2018); Harigaya and Leedom (2020).
More recently, it has been shown that a first-order phase transition relevant to the mass generation can significantly enhance ALP dark matter production, leading to novel dynamics associated with bubble expansion and axion wave generation Nakagawa et al. (2023); Lee et al. (2024). In addition to these scenarios, various other mechanisms for ALP and QCD axion dark matter production have been extensively studied for the dark matter valid in wide parameter region. Experimental efforts are also ongoing for the large discovery space; see, e.g., Ref. O’Hare (2020).
In this paper we focus on a minimal ALP model. In this model the ALP potential consists of a single cosine term:
$V=m_{\phi}^{2}f_{\phi}^{2}\bigl{(}1-\cos\bigl{(}\phi/f_{\phi}\bigr{)}\bigr{)}.$ (1)
This potential is bounded above and below, which gives an constraint on ALP dark matter from the onset of coherent oscillations, as shown in, e.g., Refs. Marsh and Yin (2021); Nakagawa et al. (2023) by taking account of constraints for the small scale structure from Lyman-$\alpha$ and Milky Way subhalo count.
Observations of small‐scale structure place a lower bound on the redshift for the ALP being the matter Sarkar et al. (2015); Corasaniti et al. (2017); Das and Nadler (2021). In particular, the abundance of Milky Way satellite galaxies sets the lower bound Das and Nadler (2021); Fujita et al. (2024)
$z_{\rm matter}\gtrsim 5.5\times 10^{6}.$ (2)
Requiring the potential hight is dominant
$(m_{\phi}f_{\phi})^{2}>\rho_{\phi},$ (3)
with the total energy density of the ALP, $\rho_{\phi}=\rho_{\rm DM}\simeq(1+z)^{3}(1.8\times 10^{-12}\,{\rm GeV})^{4}$ at $z=5.5\times 10^{6}$, yields Marsh and Yin (2021); Nakagawa et al. (2023)
$f_{\phi}\gtrsim 4\times 10^{13}\,{\rm GeV}\biggl{(}\frac{10^{-18}\,{\rm eV}}{m_{\phi}}\biggr{)},$ (4)
by assuming the ALP composed mostly by the homogeneous mode behaving like pressureless matter.
In this work, we demonstrate that the same limit applies for even more generic initial field configurations, where nonzero-momentum modes dominate the energy density. It is clear that relativistic ALP particles cannot violate limit (4), because it is obviously not the matter. The only possible loophole to violate the limit (4) may be the non-homogenous but non-relativistic particles. However, interestingly, we find through lattice simulations that when the ALP is dominated by nonrelativistic gradient modes, it nevertheless behaves like relativistic radiation if the condition (3) is violated. Furthermore, if the typical momentum is sufficiently small, the field can form domain walls once the gradient energy redshifts and becomes comparable to the potential energy. This implies that the constraint Eq.(4) is robust. Assuming the $O(1)$ anomaly coefficient to the photon, we derive a stringent bound to the minimal ALP dark matter model.
We comment on that a generic limit on the dark photon with mass given by a dark Higgs field was pointed out. This is because if the coupling is too large, the back reaction of the abundant dark photon brings the dark Higgs into symmetric phase Kitajima et al. (2025) (see also a parametric similar limit from topological defect formation Cyncynates and Weiner (2025a, b)). This paper will show a similar result to ALP from different dynamics: if the coupling $1/f_{\phi}$ is too large, the existence of too much ALP fluctuation makes the potential effectively flat.
II Relativistic axion with non-relativistic momenta
Here we perform a lattice simulation by using Cosmolattice Figueroa et al. (2021, 2023) to study the impact on the non-homogeneous mode given that the potential Eq.(1) is a cosine term, which is non-linear in $\phi$.
Let us consider
$\phi^{0}=\bar{\phi}^{0}+\delta\phi^{0},$ (5)
where $\bar{\phi}^{0}$ is the homogeneous mode, that we take $\bar{\phi}^{0}=0,\dot{\bar{\phi^{0}}}=0$, since the ALP limit on the dark matter with the dominated homogeneous mode is known. $\dot{X}\equiv dX/dt$ with $t$ being the cosmic time. $\delta\phi$ represents the fluctuation. Throughout this paper, we use superscript 00 to denote the initial condition for the simulation.
The fluctuation $\delta\phi^{0}$ satisfies
$\langle\delta\phi^{0}\delta\phi^{0}\rangle=\int^{K_{\rm UV}}_{K_{\rm IR}}\frac{dk}{k}\,{\cal P}_{\delta\phi}(k).$ (6)
Here $K_{\rm UV,IR}$ is the momentum cutoff for the spectrum we consider. We take $K_{\rm IR}\gtrsim H^{0}$, with $H$ being the Hubble parameter, because we consider the fluctuation is adiabatic for conservativeness, and we choose the uniform density slicing for the super-horizon mode. Otherwise we will have constraint from isocurvature modes.
For the reduced power spectrum, we take the form
${\cal P}_{\delta\phi}(k)=\tilde{\phi}^{2}\left(\frac{k}{m_{\phi}}\right)^{q},$ (7)
with $\tilde{\phi}$ being a normalization. It is white noise with $q=3$ and scale invariant for $q=0$. For $q\neq 0$, this gives
$\langle\delta\phi^{0}\delta\phi^{0}\rangle=\frac{\tilde{\phi}^{2}}{q}\left((\frac{K_{\rm UV}}{m_{\phi}})^{q}-(\frac{K_{\rm IR}}{m_{\phi}})^{q}\right),$ (8)
while the $q\to 0$ limit gives the case for $q=0$.
To study the late time cosmology, we solve the equation of motion,
$\ddot{\phi}-\frac{1}{a^{2}}\Delta\phi_{i}+3H\dot{\phi}=-\partial_{\phi}V_{{\rm eff}},$ (9)
by assuming a radiation dominated Universe for our purpose. $a$ is the scale factor, which is taken to be unity, $a^{0}=1$, at the initial simulation time.
In Fig.1 we show the solution for the energy densities normalized by $a^{3}$. We take $\tilde{\phi}=f_{\phi},K_{\rm UV}=m_{\phi}$ (left panel), $\tilde{\phi}=10f_{\phi},K_{\rm UV}=m_{\phi}$ (middle panel) and $\tilde{\phi}=40f_{\phi},K_{\rm UV}=1/2m_{\phi}$ (right panel). In these panels, we take $H^{0}=0.5m_{0},m_{\phi}^{2}=5m_{0}^{2},q=2,~{}{\rm and}~{}K_{\rm IR}=H_{0}$. Here $m_{0}$ is the machine unit.
Fig. 1: The (conformal) time evolution of the energy density normalized by $a^{3}$. The purple, green, blue and yellow lines denote the kinetic energy, gradient energy, potential energy, and the total energy, respectively. In all panels, we take $H^{0}=0.5m_{0},m_{\phi}^{2}=5m_{0}^{2},q=2,~{}{\rm and}~{}K_{\rm IR}=H_{0}$. In the left, middle and right panels, we choose $(\tilde{\phi},K_{\rm UV})=(f_{\phi},m_{\phi})$, $(10f_{\phi},m_{\phi})$ and $(40f_{\phi},1/2m_{\phi})$ , respecively.
Fig. 2: The histogram of $\phi/f_{\phi}$ at $\tau=8/m_{0}$.
The left panel shows the case that Eq.(3) is satisfied. In the other two panels, Eq.(3) is violated. This can be seen from the initial gradient energy density,
$\left\langle{(\nabla\phi)^{2}}\right\rangle=\tilde{\phi}^{2}\frac{\frac{K_{\rm UV}^{2+q}}{m_{\phi}^{q}}-\frac{K_{\rm IR}^{2+q}}{m_{\phi}^{q}}}{2+q},$ (10)
which is
$\left\langle{(\nabla\phi)^{2}}\right\rangle=\frac{1}{4}\frac{\tilde{\phi}^{2}K_{\rm UV}^{4}}{m_{\phi}^{2}}$ (11)
In the parameter choice for the figure, we have $\frac{1}{4}m_{\phi}^{2}f_{\phi}^{2},25m_{\phi}^{2}f_{\phi}^{2},25m_{\phi}^{2}f_{\phi}^{2}$ from the left to right panels.
We can clearly see that the total energy of the left panel behaves as matter, while the ones in the other two behave more like radiation at the beginning. The initial potential energy in the left panel scales as matter in time while the other two behaves as cosmological constant, i.e. $\left\langle{V}\right\rangle a^{3}\propto a^{3}$, at the beginning and later they behave as matter.
The difference can be understood intuitively as follows. For the relevant mode $k_{\rm IR}<k<k_{\rm UV}\leq m_{\phi}$ with large amplitudes the right-hand side of Eq.(9), which is bounded by $m_{\phi}^{2}f_{\phi}^{2}$ can be neglected. Then the gradient energy becomes comparable to the kinetic energy and ALP behaves as radiation as the massless limit. The potential energy by averaging the field fluctuations in this limit is $\left\langle{V}\right\rangle=m_{\phi}^{2}f_{\phi}^{2}$, which behaves as cosmological constant.
When the gradient/kinetic energy redshifts to the potenial energy, we have two possibilities. When the amplitude is not very large, $m_{\phi}$ is comparable to the typical momentum, one can have the ALP as the matter. This corresponds to the middle panel.
When the typical momentum becomes smaller than the mass, i.e., when the field amplitude is much larger than $f_{\phi}$, the field configurations in coordinate space begin to cross the potential barriers multiple times after the kinetic and gradient energies become comparable to the potential energy. As a result, domain walls are formed. This behavior can be observed in the right panel of Fig. 1, where the gradient energy increases.
Due to a population bias among the vacua, we expect that the domain walls will collapse within a relatively short lifetime, provided the field amplitude is not extremely larger than $f_{\phi}$. The histogram of $\phi$ corresponding to the right panel of Fig. 1 is shown in Fig. 2 at $\tau=8/m_{0}$. It clearly indicates that the field values are localized around multiple vacua $\phi=2n\pi f_{\phi}$ with $n$ being an integer, confirming that domain walls are formed by field configurations connecting these vacua.
One cannot get the matter dominant epoch if kinetic energy/gradient energy is larger than the potential energy at $z<5.5\times 10^{6}$ even if the typical momentum is smaller than the mass. This shows again Eq.(4).
III Conclusions
We have found a generic limit for the minimal ALP dark matter Eq.(4) by considering generic field configurations. This is known if the dark matter is composed of the homogeneous mode. In this paper, we showed even if it is inhomogenous, especially even when the typical momentum of the ALP is smaller than the mass, the limit holds as well. This is because if one violates the limit and requires the gradient energy of the ALP much larger than the potential energy, it behaves as radiation irrelevant to the typical momentum of the ALP particles.
The ALP is usually defined such that it couples to a pair of photons with the coupling
${\cal L}=c_{\gamma}\frac{\alpha}{4\pi}\frac{\phi}{f_{\phi}}F_{\mu\nu}\tilde{F}^{\mu\nu}\equiv\frac{1}{4}g_{\phi\gamma\gamma}\phi F_{\mu\nu}\tilde{F}^{\mu\nu}.$ (12)
We consider in the minimal setup $|c_{\gamma}|\lesssim 1$. This is because if $c_{\gamma}\gg 1$ we need mechanisms to enhance the photon coupling, e.g., coupling the ALP to many charged chiral fermions or mixing between multiple ALPs e.g. Ho et al. (2018); Cyncynates and Thompson (2023); Murai et al. (2025). [^1] An example of the other type of loophole, which is not minimal, is to have another dark matter who gives the abundance to the ALP dark matter at late time Universe without changing much the equation of state. In this case, a more conservative limit may be, $m_{\phi}^{2}f_{\phi}^{2}<\rho_{\rm DM}^{\rm galaxycenter}$, with $\rho_{\rm DM}^{\rm galaxycenter}$ being the local dark matter density which is known to be dense such as around some galaxy centers or the center of the cuspy dwarf spheroidal galaxies.
In this case, we get a generic cosmological upper limit on the ALP-photon coupling for the minimal model, which is shown in Fig. 3 (See also Marsh and Yin (2021); Nakagawa et al. (2023)). In the case the ALP is discovered in the region violating the limit, it implies that the ALP is associated with a more complicated structure, which desires further model building and understanding to the relation to the fundamental law of physics.
We also found that if the typical ALP momentum is very small, and the fluctuation is very large, the domain wall can be formed. The domain wall can be marginary long lived, and collapse due to the population bias. This will induce gravitational wave, minicluster, and primordial blackhole. A further study on producing the large low momentum mode, such as with stimulated emission Moroi and Yin (2021a, b); Yin (2023); Sakurai and Yin (2025), will be a interesting topic.
[…]
Note added While completing our study, we noticed a paper posted on arXivHarigaya et al. (2025), in which the authors discuss a similar bound.